OpenStudy (anonymous):

The simple interest formula I equals p times r times t gives the interest earned on an investment of p dollars at interest rate r for a time of t years. How much does Hector need to invest at a rate of 2% in order to earn $25 interest in 3 years?

OpenStudy (anonymous):

I=p*r*t ==> 25=p*2/100*3 => p=416,66.. usd